# Limitations of Existing Work, and Guidelines for Future Research

### v1

@@ 5.5 Different definitions for public and private sectors
After extensive debate within the fair-AI community, it is well accepted that different fairness definitions may be necessary for different decision-making contexts (Saxena et al., 2019; Mehrabi et al., 2021). This is likely true of spatial fairness definitions as well, given the diverse criteria and objectives in different decision-making domains. For instance, public policy-driven spatial decisions aim for fairness without profit, such as equitable assignment of students to schools or optimal locations for public services such as hospitals to ensure greater access. In contrast, private companies may seek profit while aiming for spatial fairness. Moreover, industries like mortgage or insurance must adhere to specific state and federal laws prohibiting discrimination.
Hence, articulating the specific domain(s) (public or private) and industries targeted by a proposed spatial definition or method and elucidating the conditions and laws it aims to abide by upon its introduction would facilitate subsequent research endeavors. This includes the evaluation of whether the stipulated conditions and laws are actually adhered to in practice. Since a particular spatial fairness definition or technique may not be generalizable to all spatial decision-making scenarios, doing so would establish a foundation for practical applications and simultaneously ground the work in the possibility that the proposed method is not guaranteed to be generalizable beyond the specified contexts.
5.6 Serve as a tool to measure spatial inequities
Ideally, a spatial fairness definition or methodology should also serve as a tool to detect (or audit for) the presence of spatial inequities. This is not always the case, as Sacharidis et al. (2023) show the measure proposed in Xie et al. (2022) cannot be meaningfully used to audit for spatial fairness or to find which regions are spatially unfair. They go on to propose a methodology that can be utilized to audit for spatial fairness with respect to positive/negative outcomes, but do not consider the spatial regions’ correlation with protected characteristics. Fortunately, there is some work in the social sciences exploring auditing for inequities, including spatial inequities, which may serve as a foundation for spatial fairness and auditing. For example, Saporito and Maliniak (2022) propose a technique to distinguish between legislative districts that are “unnaturally packed” with racial minorities (i.e., are segregated) from districts that are not to aid the study of racial gerrymandering.
Tools for auditing spatial fairness can have significant, meaningful impact in multiple ways. First, by helping computationally implement recent initiatives passed by the justice department to audit lending practices to assess when an individual’s race, national origin or neighborhood had an impact in their access to credit. Second, by helping automate spatial auditing analyses of changes in compositions of redlined neighborhoods over time to support appropriate remedial policy decisions. For example, Perry and Harshbarger (2019) found that housing proposals by some 2016 presidential hopefuls meant to address persisting harms from redlining would not have benefited African Americans as much as intended because the racial makeup of redlined neighborhoods in some areas had changed over time. Spatial fairness tools can make such auditing easier to carry out, helping ensure policies benefit the intended recipients.
5.7 Good faith effort to explore relevant environmental, social, and economic issues in social planning
Fair-AI researchers’ familiarity with public policy and economic development recommendations regarding location (Squires & Kubrin, 2005; Lewis & Hamilton, 2011; Li et al., 2013; Kaplan, 1998; Osypuk & Acevedo-Garcia, 2010) would significantly help integrate their research into practical applications. While it cannot be expected for all computer science research to abide by public policy guidelines, fair-AI researchers should strive to be familiar with public policy and economic development recommendations that are directly relevant to their area of research. Integration of practices recommended to reduce inequities would ensure fair-AI techniques do not unintentionally perpetuate such inequities themselves if implemented in the real world.
@@ 6 Conclusion
This work argues for the importance of spatial fairness. We delineate the need to address persisting bias in spatial decision-making contexts and legal efforts proposed to address it, and outline problems unique to spatial analysis that compound the difficulties in addressing spatial bias computationally. Furthermore, we identify drawbacks of the handful of spatial fairness definitions and techniques proposed so far to highlight why they are insufficient. Finally, we establish a set of guidelines to support future spatial fairness research avoid similar pitfalls and achieve maximal impact with adoption in the real world.
Impact Statement
This position paper presents an argument (or position) to bring attention to spatial fairness in the field of machine learning with the goal of motivating future work on mitigating location-related biases in society. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v2

@@ 5.2.1 Tailoring Fairness to Domain and the Law
Moreover, extensive debate in the fair-AI community has established that no single fairness definition is appropriate for all decision-making contexts (Saxena et al., 2019; Mehrabi et al., 2021). The same is likely true for spatial fairness, given that decision-making objectives and acceptable trade-offs vary widely across domains. For example, spatial decisions in the public sector are often driven by equity-oriented goals rather than profit, such as an equitable assignment of students to public schools. In contrast, private sector decision-making typically involves profit incentives alongside fairness objectives, while regulated industries such as mortgage lending and insurance must also satisfy specific federal and state anti-discrimination requirements.
Consequently, spatial fairness definitions should clearly specify the intended domain, industry context, and the legal and policy constraints they aim to satisfy. This enables meaningful compliance evaluation and delimits the scope of generalizability, recognizing that a definition suitable for one spatial setting may not transfer to others. Explicitly articulating these assumptions establishes a foundation for deployment and rigorous, comparable future research.
5.2.2 A Joint Protected Attribute Framework
Treat location as a protected or immutable characteristic. Immutability is deeply associated with anti-discrimination law in the U.S. (Hoffman, 2010). Recent interpretations are a characteristic “that either is beyond the power of an individual to change,” [In re Acosta, 19 I. & N. Dec. 211, 233-34 (B.I.A. 1985)] or in other terms, “either unchangeable in absolute terms or so fundamental to identity or conscience that individuals effectively cannot and should not be required to change it” [In re Acosta, 19 I. & N. Dec. 211, 233-34 (B.I.A. 1985)] or “changing it would require great physical difficulty, such as requiring a major physical change or a traumatic change of identity.” [Watkins v. U.S. Army, 875 F.2d 699, 726 (9th Cir. 1989)] (Hoffman, 2010), thereby establishing a characteristic need not be unchangeable to be considered immutable under anti-discrimination law. This, in addition to ongoing efforts to afford protection against discrimination on the basis of zip codes, brings us to the next guideline: consider location a protected attribute along with traditional legally protected attributes like race, or at least an immutable one. We outline example use cases below.
In domains with a demonstrated history of location-related discrimination, zip codes could be considered a joint protected attribute with the legally protected attributes it perpetuates bias against. For example, an intuitive measure of individual fairness is to treat similar people similarly (Dwork et al., 2012). With location and race as a joint protected attribute, two applicants, $i$ and $j$, who are similar to each other on non-protected aspects should receive similar decisions, $D(i)\approx D(j)$, regardless of their race and location. When impractical, location should be treated as an immutable attribute. For example, some fair- and explainable-AI research aims to provide applicants with actionable points to increase their likelihood of a positive decision (Ustun et al., 2019). Suggested actions should, however, be reasonably achievable. For example, a model may recommend increasing annual salary by \$2,000 to improve odds of loan approval, but if the salary of an applicant with a PhD degree is low for someone with a PhD degree, the model cannot recommend they undo their PhD. Similarly, residence should be treated as immutable for low-income families, i.e., something a low socio-economic status applicant cannot reasonably change since they commonly do not have the freedom to make this choice, sometimes even if they can afford to (e.g., due to source of income discrimination (see Section 2.1)).
5.2.3 Disrupt Feedback Loops via Exploration
Incentivize exploration to break cycles of neglect. To prevent the entrenchment of historical biases, models should encourage an exploration-exploitation trade-off. By incentivizing collection of fresh data from under-sampled regions rather than solely optimizing for known rewards, systems can disrupt feedback loops where marginalized neighborhoods are systematically neglected or unfairly targeted based on a prejudiced past. This approach draws on research into feedback loops in predictive policing models (Ensign et al., 2018), and work exploring how to break such cycles (Joseph et al., 2016; Jabbari et al., 2017).
5.2.4 Mitigating MAUP Resistance
Ensure robustness to the Modifiable Areal Unit Problem (MAUP). Since results in spatial analyses can vary substantially depending on how geographic units are defined, spatially fair systems should be robust to MAUP. MAUP manifests through two primary mechanisms: the zonal effect, in which results change when geographic boundaries are redrawn, and the scale effect, in which outcomes shift as data are aggregated to coarser or finer spatial resolutions. These sensitivities can undermine fairness claims. For example, a system may appear equitable under one administrative partition (e.g., census tracts) but exhibit disparate impact under another (e.g., zip codes), or fairness assessments may change simply due to arbitrary aggregation choices.
To mitigate these risks, spatial fairness research should incorporate robustness checks across multiple (plausible) zoning schemes and spatial scales, and explicitly report how fairness metrics vary under alternative unit definitions. Related work in spatial statistics and geography offers ideas for mitigating MAUP in applied settings, such as traffic safety (Xu et al., 2018) and spatial public health analyses (Tuson et al., 2019; Tuson, 2022; Hennerdal and Nielsen, 2017). Other work explores the design of zone-based districts for constructing partitions that align with domain objectives while avoiding MAUP-driven artifacts (Martin, 2003; Tuson, 2022). Spatial fairness research can build on these foundations and domain expertise to determine the appropriate spatial resolution and partitioning scheme for the problem at hand, and ensure fairness evaluations remain well-founded when assessed under plausible alternative representations.
5.3 Validation and Monitoring
5.3.1 Empirically Verifying Bias Mitigation
Explicitly measure bias reduction against protected attributes. Although spatial fairness research is motivated by location’s correlation with protected attributes, recent works rarely demonstrate that their methods actually mitigate bias against legally protected attributes (Shaham et al., 2022; Sacharidis et al., 2023). To ensure real-world utility, future work must explicitly evaluate spatial bias with respect to protected attributes. Researchers must compare bias levels before and after applying their methodology, while also reporting trade-offs with other objectives (e.g., prediction accuracy). For example, if a bias mitigation intervention transforms spatial data $L(x)$, where $x$ is a protected attribute, into $L^{\prime}(x)$, it is essential to quantify how much this actually alleviates bias against the protected attribute rather than simply assuming the spatial correction is sufficient.
5.3.2 Continuous Auditing and Monitoring
Establish ongoing compliance and monitoring pipelines. Models deployed in sensitive or high-impact domains should be accompanied by an ongoing compliance and monitoring pipeline. Since demographic and spatial patterns can change over time (e.g., through gentrification, migration, or infrastructure investments), models that were initially evaluated as spatially fair may later exhibit different disparities. Regular re-auditing is therefore essential to ensure continued alignment with legal and policy requirements, and to detect emerging spatial biases before they can transform into harmful outcomes. Where practically plausible, monitoring should include periodic disparate impact assessments with clearly defined triggers for human review.
6 Alternative Views
We address three credible counterarguments to our position.
Spatial inequity is outside the scope of algorithm design. One view is that spatial inequities are fundamentally structural, and therefore should be addressed through policy and institutional reform, rather than technical modifications to models. Critics might argue that “fixing” spatial bias in a mortgage or ride-hailing model is a band-aid that distracts from the root cause. We agree that algorithmic interventions are not a substitute for policy reform. However, data-driven systems are no longer passive observers of the city, they actively reshape it. Decision-making systems that optimize for short-term efficiency on historically biased data can entrench and accelerate spatial disadvantage, effectively “automating” the legacy of practices like redlining. Technical spatial fairness is therefore a necessary containment mechanism, a guardrail to prevents AI from amplifying structural inequities while longer-term policy reforms take hold.
Existing fair-AI approaches are sufficient. Another view is that spatial bias does not warrant dedicated attention because it can be handled with existing fair-AI approaches. We argue that this view ignores the unique statistical properties of spatial data that undermine standard assumptions in much fair-AI work. this conclusion is premature. As detailed in Section 3, spatial data challenges that violates the independence assumption. Furthermore, standard fairness metrics are brittle to MAUP. Without explicit handling of scale and zoning effects, standard fairness constraints are mathematically insufficient for spatial applications.
The efficiency-vs-fairness argument. Finally, practitioners in high-stakes domains may argue that imposing spatial fairness constraints will degrade system utility, increasing costs or response latencies, thereby harming the very populations the system aims to serve. We acknowledge there is often a short-term trade-off between fairness and optimization. However, ignoring spatial fairness incurs significant long-term costs. Spatially biased systems risk regulatory backlash (e.g., the legislative proposals in Illinois) and loss of user trust. Moreover, incorporating exploration into spatial models can disrupt feedback loops, potentially discovering untapped demand in underserved neighborhoods and leading to a more robust, globally optimal system rather than one stuck in a local optimum based on historical prejudice.
@@ 7 Conclusion
This work argues for the importance of spatial fairness. We delineate the need to address persisting bias in spatial decision-making contexts and legal efforts proposed to address it, and outline problems unique to spatial analysis that compound the difficulties in addressing spatial bias computationally. Moreover, we identify drawbacks of the handful of spatial fairness work proposed so far to highlight why they fall short. Finally, we establish a set of guidelines to support future spatial fairness research avoid similar pitfalls and achieve maximal impact with adoption in the real world.